OpenStudy (anonymous):

limt x-->inf x/x+5 using divergent test limt x-->inf x/x+5 using divergent test @Mathematics

myininaya (myininaya):

\[\lim_{x \rightarrow \infty}\frac{x}{x+5} \cdot \frac{\frac{1}{x}}{\frac{1}{x}}=\lim_{x \rightarrow \infty}\frac{1}{1+\frac{5}{x}}=\frac{1}{1+0}=1\]

OpenStudy (anonymous):

ohh wowww...waht an idiot